OpenStudy (anonymous):

what is theta if my complex root is z=-5i?? i keep getting an error on my calculator

OpenStudy (psymon):

Well, you should get undefined :P

OpenStudy (psymon):

BUT

OpenStudy (psymon):

The conversion to get theta is \[\tan \theta = \frac{ b }{ a }\] In this case, b is -5 and a is 0. So yeah, you get: \[\tan \theta = \frac{ -5 }{ 0 }\] Well what that tells me is I want to know where tangent is undefined. So what angle of tangent gets me an undefined answer?

OpenStudy (anonymous):

i know but the problem asks for a complex fifth root -_- :/

OpenStudy (anonymous):

thank you for that explanation.. how would you solve this to get Zo?? that's all i need to solve this

OpenStudy (psymon):

I was going to work up to that, sorry if you didn't need any of the previous stuff. Well, if you want a root you need this: \[\sqrt[n]{r}(\cos \frac{ \theta }{ n }+isin \frac{ \theta }{ n })\] So you shouldn't get an error, you should just take the 5th root of r and then divide the angle by 5.

OpenStudy (psymon):

And if I frustrate you then just ask me to leave, I'm not out to annoy people on here.

OpenStudy (anonymous):

but isn't r in that equation \[\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

oh sorry no you're not frustrating at all!

OpenStudy (psymon):

That gets you r at first, yes. But once you get r, you then take the 5th root of it.

OpenStudy (anonymous):

I'm frustrated with my brain and the problem -__-

OpenStudy (anonymous):

I'm not frustrated lol sorry again

OpenStudy (psymon):

So assuming you've made it far enough to get r and theta, we would have: \[5(\cos \frac{ 3\pi }{ 2 }+isin \frac{ 3\pi }{ 2 })\] Is this what you got before worrying about the roots?

OpenStudy (anonymous):

yup. I actually got stuck when i tried to find theta.. i forgot this formula. Thank you for reminding me :)

OpenStudy (psymon):

Oh xD Lol, gotcha. Well now you just need to take the 5th root of r, which is 5, and then divide the angles by 5. Do you need all 5 complex roots?

OpenStudy (anonymous):

yes, i need to answer in standard form from Zo-Z4. If you want to provide answers that's kind of you but I will only use them to check my work :) that's such a relief! thank you!!

OpenStudy (psymon):

Well, given the formula, what'd we get for the 1st root?

OpenStudy (anonymous):

one quick question..the formula in my head is \[Zo=\sqrt[5]5{?}(\cos \frac{ \theta+2\pi(0) }{ n })+isin (\frac{ \theta+2\pi (0) }{ n })\] is that correct? your formula looks a bit different

OpenStudy (psymon):

Well I didn't throw in the next roots, but yes, that looks fine. So you would essentially divide 2pi by 5 and then add that to your current root until you gt all 5 you need. So your first root should have had 3pi/10, correct?

OpenStudy (anonymous):

sorry one sec.. i'm rechecking my calculations :)

OpenStudy (psymon):

no worries.

OpenStudy (anonymous):

ok my answer doesn't add up. I was taught to calculate theta by tan(b/a) and that gives my undefined. so i'm back to square one :/ I'm sorry, can you explain another way i can calculate theta?

OpenStudy (psymon):

Right, it should be undefined. But there are two angles where tangent is SUPPOSED to be undefined. Do you know those two angles?

OpenStudy (anonymous):

pi/2 and 3pi/2 oh my.. you are a life save :O i didn't even see that!!!!!! and i have the table memorized :O

OpenStudy (psymon):

Lol, yeah xD Now since our rectangular point was -5, this means we are down 5 and thus at 3pi/2. Now you can continue :P

OpenStudy (anonymous):

so because the 5 is negative.. only 3pi/2 applies? i've never seen that before *_*

OpenStudy (anonymous):

OK! i think i have my first answer! 3pi/10? and i just add 3pi/10 to each argument after that. correct?

OpenStudy (psymon):

You don't add 3pi/10, you add 2pi/n. So you would add 2pi/5

OpenStudy (anonymous):

7pi/10 is my next theta correct? before i add it to 2pi and divide by 5? i hope i'm not confusing you

OpenStudy (psymon):

Right, 3pi/10 is the first, 7pi/10 the 2nd, 11pi/10 the first, etc xD

OpenStudy (anonymous):

you are seriously amazing. i wish you were my teacher :)

OpenStudy (anonymous):

Thank you to unimaginable ends!!

OpenStudy (psymon):

Nah, Im not nearly as good as most people on here, lol.

OpenStudy (psymon):

I think I'm one of the least knowledgable to be honest, lol. People make me feel stupid all thetime x_x

OpenStudy (anonymous):

oh no no no no. please allow yourself to receive this one complement :) you really helped me sort out a 2 hour problem!!!! I calculated Z1= cos (27/50)+iSin (27/50) in standard form. does that look correct from your work? :)

OpenStudy (psymon):

Where is this coming from? Just curious since we lost our r and our pi and all of that, lol.

OpenStudy (anonymous):

because i am now adding theta to 2pi(K) all over 5. when i added Cos 7pi/10 to 2pi(K=1) all over 5 i got 27pi/50.. i knew it looked wrong

OpenStudy (psymon):

So you accidentally left out pi then. That and no idea where your radius went. Okay, so this is what we have: \[\sqrt[5]{5}(\cos (\frac{ 3\pi }{ 2 }+\frac{ 2k \pi }{ 5 })+isin(\frac{ 3\pi }{ 2 }+\frac{ 2k \pi }{ 5 })\] Sorry bout thewait. So know you just make k be 0, 1, 2, 3, 4.

OpenStudy (psymon):

Oops xD, should have divided the 2's by 5 as well.

OpenStudy (anonymous):

That's ok! i can follow! i just get confused when k=1. that means i have to add 7pi/10 to 2pi and put ALL of that over 5. correct? that gives me 27pi/50

OpenStudy (psymon):

No, you keep constantly adding 4pi/10 each time. So you start with 3pi/10, which is your first. 2nd = 3pi/10 + 4pi/10 = 7pi/10 3rd = 7pi/10 + 4pi/10 = 11pi/10 4th = 11pi/10 + 4pi/10 = 15pi/10 5th = 15pi/10 + 4pi/10 = 19pi/10 And those are your answers,

OpenStudy (anonymous):

so i don't add the 2pi(k) to my new argument? sorry if i'm tiring you, i'm trying to understand this :S

OpenStudy (psymon):

No, you only had 2kpi/5 and you just keep constantly adding that same amount.

OpenStudy (anonymous):

so my new answers in standard form look like: \[z1=(\cos7\pi/10)+(isin7\pi/10)\] and \[Z2: (\cos 11\pi/10)+(isin 11\pi/10)\]

OpenStudy (anonymous):

correct?

OpenStudy (psymon):

You keep dropping your r in front xD

OpenStudy (anonymous):

oh add that in the front mentally lol sorry! my brain is understanding stuff XD

OpenStudy (psymon):

Yeah, youre doing it right justcant drop the 5th rt(5) xD

OpenStudy (anonymous):

thank you!! i understand it now :)

OpenStudy (psymon):

Glad to hear it ^_^

OpenStudy (anonymous):

:))))